DIRC Operations Weekly Report
Period: Saturday 2006/06/03 - Friday 2006/06/09
Oncall: Emmanuel Latour, Rolf Andreassen
Fast Monitoring: Emmanuel Latour, Rolf Andreassen
File location: $BFROOT/www/Detector/DIRC/WeeklyReports/report_20060609.html
Plot directory: $BFROOT/www/Detector/DIRC/WeeklyReports/Plots/report_20060609 (empty)
PEP peak lumi: 10.3*1033
(previous week: 10.4 1033)
BaBar integrated lumi: 2.943 fb-1
(previous week: 2.669 fb-1)
Run 5b PEP delivery: 55.53 fb-1
Run 5 total delivery: 113.97 fb-1 Run 5 total BaBar recorded: 109.72 fb-1 including 6.26 fb-1 off-peak data
PEP activities this week:
Tue morning: characterisation of the machine
Wed: PEP PR02 micro problem+LER region 4 vacuum problem: had to remove a chamber and braise it. Beams back on Saturday (June, 10th) owl shift.
Thu: a klystron failed during swing shift due to a water leak. They replaced it.
PEP activities planned next week :
Carry on offpeak data taking
Foreseen work opportunities (PEP down periods): None.
BaBar activities/issues this week: UPS system tripped on Saturday owl shift (at about 4:30pm). Lasted 11 hours.
Loss of power in almost all EH.
Loss of monitoring/computing
Pump for cooling water turned off
Failuer of SVT chillers
UPS system finally bypassed (brought back within BaBar on Thursday while magnet was ramped down)
See DIRC logbook for details
Downtime due to the DIRC: None.
- Sat: UPS outage at BaBar.
- FEE, HV, gas system, Epics, degasing pump all off.
- Paged Matt and Jerry for the gas system. Matt came to disable the valve
- Jerry came to re-enable it.
- 2 HV boards in crate 3, sector 4, died: one was tripped and one was at 0. We replaced the 2 boards. We took a standalone run, and checked data on DQM plots for cosmics. It looked fine.
- Sun: Calibration was taken later in the night. Data was bad for sector 4, in the HV groups powered by the 2 new boards. The DQM in the morning sent an email for a bin that was high in the DIRC plots. Nicolas figured out we had not set the nominal voltages after having changed the boards. He phoned Emmanuele who did it. It affected 4 runs, but the values of high voltage were not too different, so data was finally OK.
Post processing of fast Mon had crashed. Jim Hamilton and Matthias Wittgen fixed it (the machine did not restart after the UPS outage)
- Wed: Dirc monthly report at the Ops meeting (Emmanuele)
Background summary plot
- DIRC biweekly meeting on wednesday, June 14th 2006. In particular, rehearsal for the talk at Montreal.